During the investigation we found the INSERTION LOSS HIGH on IU19 DAS board and SPAN_LOSS_HIGH alarm on IU121 CRPC board.
Suggestion: check the fiber/ the pigtail and the board is working properly.
- fibers checked at DCM board. No problem detected
- optical power check at pins RTDC and RRDC of DAS board slot 19.
- CRCP slot 121 replaced with a new one. Slight improvement of optical power received by DAS slot 19 but the alarm persist.
Make sure that the gain of the OA board is adjusted to the maximum value, which exceeds the permitted range.
To clear the insertion_loss_high alarm, the output optical power of the CRPC board in slot 121 is increased from 24.7 to 25.2, and the gain of the DAS1 board is reduced to 26db. So the optical power had been optimized.
The gain set for the optical amplifier board exceeds the maximum permitted actual gain.
The optical power had been optimize on amplifier.
The optical power had been adjust on the output of CRPC board and the gain on DAS was decreased.
Check the ALC link status to be sure that attenuation is correct. The ALC alarm threshold is 2dB.
***as an example, be aware that attenuation should be lower than node gain.